OpenStudy (anonymous):

Find the quadratic equation with solutions 3 and -3/7

OpenStudy (anonymous):

(x-3)(x+3/7 )

OpenStudy (mertsj):

(x-3)(x+3/7)=0

OpenStudy (mertsj):

x^2 -18/7x -9/7=0

OpenStudy (mertsj):

7x^2 - 18x -9 = 0